OpenStudy (hang254):

A kindergarten school has square mats in two colors, red and blue. The ratio of the perimeter of the red mat to the perimeter of the blue mat is 1:2. The area of the blue mat is 100 square feet. What is the area of the red mat? 25 square feet 75 square feet 10 square feet 50 square feet

OpenStudy (valpey):

Twice the perimeter is four times the area. Or conversely, half the perimeter is one quarter the area.

OpenStudy (hang254):

so, since 10(10)=100, it is 10?

OpenStudy (anonymous):

First convert the area into perimeter. The blue mat has 100 square feet. Therefore, each side must have 10 ft since it's square. Then, 10 x 4 since square has four sides. so you get 10 x 4 = 40 ft for the blue mat. The ratio is 1:2 (red:blue) 1:2 = x: 40 so, change it into fraction 1/2 = x/40 x = (1/2) x40 = 20ft remember the question is asking area and 20ft consists of four sides since it's perimeter. Therefore, divide 20 with 4. 20/4 = 5ft Red mat's each side has 5ft. Therefore, the area has to be 25 square meter. Hope it helped.. =)

OpenStudy (hang254):

Thank you :)